# Splat Feature Solver

### v1

@@ Conclusion
Appendix A Tikhonov Guidance
@@
Appendix B Post Lifting Aggregation Details
@@
We use the cuML implementation of HDBSCAN for clustering, and the initial results are generally reliable. However, as noted in the “Tikhonov Guidance” section of the appendix, small noise in the observations—manifesting as nearly identical rows—can produce large errors. A straightforward remedy is to identify and discard those unreliable observations.
@@
Why can we safely discard them? Recall our problem definition: “each primitive should admit a single descriptor.” In other words, nearly identical rows should yield nearly identical features. When they do not, the discrepancy must arise either from the solver or from corrupted observations. Because our subsequent queries compare cluster-level attention scores, noise that affects all clusters equally (cluster-independent noise) merely shifts every cluster’s score by a constant and does not alter their relative ordering.
@@
By contrast, cluster-dependent noise—often introduced by nearby objects, as illustrated in Fig. 9—does change those relative scores. After clustering, we therefore generate a pseudo-mask for each cluster and compute its mIoU against the original masks, filtering out any clusters whose pseudo-mask indicates unreliability.
Appendix C Training Pipeline Setting
@@
Appendix D Lifting Implementation
@@
Appendix F Property Justification
@@
Here we substitute the residual $\epsilon$ back to the Equ.30:
@@
When we expand the above equation, we get the expectation of the gradient is the Equ.34
@@
And apparently, we can get everything canceled out, except this term shown in Equ.35
@@
To experimentally validate our theoretical prediction, we conducted the following experiment. We trained both 3D Gaussian Splats and 2D Gaussian Splats separately using random backgrounds, then computed the mean and variance of their final alpha distributions. If the mean exceeds a threshold $\psi$ = 99.6, our theory is confirmed. We selected $\psi$ = 99.6 because color values are quantized into 256 levels, and additional floating-point precision does not improve RGB rendering. For comparison, we also ran the 3D Gaussian Splats experiment with a black background. As shown in Tables 9 and 10, the results align well with our theoretical proof.
@@
Appendix G Feature Visualization
Details of the feature visualization, one can check Fig.3, Fig.4 and Fig.5. For multiple kernel feature visualization, one can check Fig.10. There are plenty of the content in the appended video. Please feel free to check
Appendix H Threshold

### v2

@@ 6 Conclusion
Appendix A Multimedia Supplementary
In the supplementary we submitted, we have additional embedding website. The website is local, so one needs to open it by uncompressing the website package and click the index.html. We have additional code zip file with detail README guidance. To preview the overall result, one can ckeck our video.
Appendix B Tikhonov Guidance
@@
Appendix C Denoising Properties of Post-Lifting Aggregation
The Post-Lifting Aggregation module is a critical component for denoising input errors. It is effective not only at resolving inconsistent segmentation masks but also at mitigating errors introduced by noisy camera poses.
Implementation
As described in the main paper, we first cluster the initial lifted features using HDBSCAN. Next, we project each 3D cluster onto all training image frustums to identify reliable 2D masks that overlap with the cluster. Masks with an Intersection-over-Union (IoU) overlap above 75% are deemed reliable, while others are discarded during feature aggregation. Finally, we aggregate features only from these reliable masks—merging them into a single feature representation—and assign this unified feature to all splats within the corresponding cluster.
Intuition
@@ Appendix C Denoising Properties of Post-Lifting Aggregation
As noted in the ”Tikhonov Guidance” section (Appendix B), small noise in observations—manifesting as nearly identical rows in the system matrix—can lead to large errors. A straightforward remedy is to identify and discard these unreliable observations.
@@
Why is it safe to discard them? Recall our problem definition: ”each primitive should admit a single descriptor.” Theoretically, nearly identical rows (rays hitting the same primitive) should yield nearly identical features. Discrepancies imply errors either from the solver or corrupted observations. Since our queries rely on cluster-level attention scores, noise that affects a cluster uniformly (cluster-independent noise) merely shifts the score by a constant without altering the relative ordering.
@@
In contrast, cluster-dependent noise—often introduced by occlusions or semantic bleed from nearby objects (see Fig. 7)—distorts these scores. By generating a pseudo-mask for each cluster and computing its IoU against the original input masks, we can effectively filter out clusters where the pseudo-mask indicates unreliability.
Crucially, this mechanism makes our method extremely robust to noisy camera poses. Misaligned poses result in low overlap between the projected cluster and the 2D mask, causing those corrupted observations to be automatically filtered out. We validate this robustness in the following experiments.
Experiments
We first demonstrate robustness against inconsistent masks in Tab. LABEL:tab:ablation and Tab. 4. Our aggregation module yields a direct gain of 1.9% mIoU even with accurate camera poses, simply by resolving mask inconsistencies.
To evaluate robustness against camera pose noise, we conducted a controlled experiment where we artificially corrupted the poses of a random subset ($N$) of every 100 images in the training set. The splat geometry was fixed (trained on clean poses). The noise injection was defined as follows:
  • Rotation: We added Gaussian noise with $\sigma=1^{\circ}$ to roll, pitch, and yaw.
  • Translation: We added Gaussian noise with $\sigma=\frac{\text{scene scale}}{100}$.
We then applied our solver to this noisy data. The quantitative results, presented in Tab. 3, demonstrate that our method maintains high performance even when a significant fraction of poses are corrupted. Additionally, we provide a qualitative comparison between results obtained from clean poses versus those with 50% pose corruption in Fig. 9.
Figure 9: Qualitative robustness against noisy camera poses. The columns display (from left to right): the input RGB images, semantic segmentation results using clean camera poses, results with 50% of camera poses corrupted (50/100), and the Ground Truth (GT) segmentation masks. Although there is a slight drop in visual quality, our method maintains consistent segmentation performance even under severe pose noise.
Table 3: Robustness against noisy camera poses. The first column indicates the number of images corrupted (per 100 images). The remaining columns report the downstream mIoU scores. Note that cosine similarity is not used here because corrupted camera poses invalidate the ground truth projection alignment. Our method maintains high performance ($\approx 65\%$ mIoU) even when up to 20% of camera poses are noisy, demonstrating significant robustness.
Corrupted Images (per 100) Figurines Ramen Teatime Waldo Kitchen Mean
100/100 1.43 56.0 18.1 36.8 28.1
50/100 64.7 71.7 58.9 48.1 60.8
20/100 72.1 63.3 69.1 55.2 65.0
10/100 68.8 64.8 63.9 57.4 63.7
3/100 65.6 64.2 63.4 59.5 63.2
2/100 66.9 64.2 63.4 59.1 63.4
1/100 66.2 63.0 64.7 59.8 63.6
[…]
Table 4: Mean IoU (mIoU%) on LeRF OVS for various methods. These are additional experimental results using our data for the ablation study. We use different backbones to train the splats and evaluate the resulting mIoUs
Method Figurines Ramen Teatime Waldo Kitchen Means
DrSplat (3D Query) 47.48 36.66 66.16 47.48 49.45
Gsplat (3DGS) W/o P W/o T (Threshold = 0.65) 55.30 47.81 63.48 49.78 54.09
Gsplat 2DGS Backbone W/o P W/o (Threshold = 0.65) 62.00 56.04 66.34 51.07 58.86
Gsplat (3DGS) W/P w/o T (Threshold = 0.65) 56.55 62.24 68.04 53.01 59.96
Gsplat (2DGS) W/P W/o T (Threshold = 0.65) 67.83 60.20 66.96 47.44 60.62
Inria Trained Result W/o P, Naive (W/o Tikhonov) (Threshold) 60.06 55.30 68.33 47.73 57.85
Inria Trained Result W/o P, (Threshold=0.65) 61.30 54.20 67.60 45.01 57.03
Inria Trained W/p W/ Tikhonov (1.2 / 1.2) 61.70 53.75 67.80 49.60 58.21
Inria Trained Result W/p w/ (Tikhonov square, no sigmoid) 71.96 58.16 69.29 52.73 63.03
Inria Trained Result W/p (Threshold 0.65) + W/T 64.76 61.70 71.63 54.71 63.20
Inria Trained W/p W/o Tikhonov Square, w/o sigmoid 65.45 58.55 72.04 50.38 61.61
Inria Trained W/p W/o (Tikhonov $\lambda$ w/o squeer) W/ hist selection 72.19 64.28 66.08 56.69 64.81
Inria Trained W/o T W/p W/ Dynamic Threshold 69.42 63.89 69.53 61.55 66.10
Inria Trained W/ T W/ P W/ Dynamic Threshold 67.64 62.34 68.48 62.11 65.14
Table 5: Effect of Tikhonov Guidance ($\lambda_{\rm square}/\lambda_{\rm reg}$) on mIoU (%). $\lambda$ Equ denotes the $\lambda$ parameters used when solving the Feature Lifting Equation. When using a text query to generate the attention map, we apply $\lambda$ Proj to obtain the 2D attention map. The experiments show that setting both $\lambda$ values to 1.2 yields optimal results.
$\lambda$ Equ / $\lambda$ Proj Figurines Ramen Teatime Waldo Kitchen Means
0.5 / 1.0 59.93 53.77 66.43 41.32 55.36
0.8 / 1.0 62.77 54.02 67.38 43.90 57.02
1.0 / 1.0 61.30 54.20 67.60 45.01 57.03
1.1 / 1.1 61.59 53.96 68.93 47.72 58.05
1.2 / 1.2 61.70 53.75 67.80 49.60 58.21
1.35 / 1.35 61.80 53.44 65.56 50.36 57.79
1.5 / 1.5 61.80 53.40 65.53 50.40 57.78
Table 6: Our methods generated masks versus LAGA generated mask on 3D-OVS dataset, as we mentioned in Fig.6, room scene is not labeled correctly in ground truth, therefore, we provide two results on room scene. The former one is the result measured on original 3DOVS dataset, and the left one is the fixed results. In both comparison, we are at least the same compare to the current state of the art method.
Methods Bed Bench Lawn Room Sofa Means
LAGA 84.9 83.3 93.6 93.1/86.2 76.5 86.3/84.9
Ours 90.2 93.2 89.0 85.5/92.7 73.6 86.3/87.7
Table 7: Weight Summation of each rows statistic property. In the Property.2, we assume that the row summation is 1, and in the main thesis, we state that this property usually holds in the splats rendering. Here is experimental results regarding the final weight summation of each pixel. The number is with a %. The dataset we are using is LeRF
Method Metrics Figurines Ramen Teatime Waldo Kitchen Means
3DGS(means) Black Background 99.49 97.92 99.40 99.86 99.17
3DGS(Std) Black Background 1.37 1.54 0.67 0.18 0.94
3DGS(means) 99.74 99.85 99.83 99.91 99.17
3DGS(Std) 0.01 0.04 0.02 0.05 0.04
2DGS(means) 99.86 99.46 99.83 99.91 99.77
2DGS(Std) 0.04 0.45 0.05 0.05 0.15
Table 8: Weight Summation of each rows statistic property. In the Property.2, we assume that the row summation is 1, and in the main thesis, we state that this property usually holds in the splats rendering. Here is experimental results regarding the final weight summation of each pixel. The number is with a %. The dataset we are using is 3DOVS
Methods Bed Bench Lawn Room Sofa Means
3DGS(means) 99.92 99.80 99.67 99.64 99.93 99.79
3DGS(Std) 0.01 0.04 0.08 0.04 0.01 0.04
2DGS(means) 99.92 99.93 99.95 99.92 99.92 99.93
2DGS(Std) 0.01 0.01 0.00 0.01 0.01 0.01
Appendix D Training Pipeline Setting
@@
Appendix E Lifting Implementation
@@
Appendix G Property Justification
@@
Here we substitute the residual $\epsilon$ back to the Equ.24:
@@
When we expand the above equation, we get the expectation of the gradient is the Equ.28
@@
And apparently, we can get everything canceled out, except this term shown in Equ.29
@@
To experimentally validate our theoretical prediction, we conducted the following experiment. We trained both 3D Gaussian Splats and 2D Gaussian Splats separately using random backgrounds, then computed the mean and variance of their final alpha distributions. If the mean exceeds a threshold $\psi$ = 99.6, our theory is confirmed. We selected $\psi$ = 99.6 because color values are quantized into 256 levels, and additional floating-point precision does not improve RGB rendering. For comparison, we also ran the 3D Gaussian Splats experiment with a black background. As shown in Tables 7 and 8, the results align well with our theoretical proof.
@@
Appendix H Feature Agnostic and Kernel Agnostic
Feature Agnosticism
Recall from Section 3.2 that our solver is proven to be optimal under any convex loss function. Theoretically, this guarantees a tighter error bound compared to heuristic or optimization-based approaches such as Chacko et al. (2025); Jun-Seong et al. (2025); Guo et al. (2024); Joseph et al. (2024). While Table 1 demonstrated that our method consistently achieves a cosine similarity above 80%, that result alone did not quantify our superiority over specific baselines.
To address this, we provide a detailed comparison against current feature-agnostic methods, specifically Guo et al. (2024); Chacko et al. (2025). We exclude Dr.Splat Jun-Seong et al. (2025) from this comparison because its pipeline necessitates compressing original features (e.g., 512 dimensions) down to low-dimensional subspaces (e.g., 32 dimensions), thereby losing feature generality. Similarly, grouping-based methods are excluded as they cannot operate without mask guidance. As shown in Table 9, the high similarity range (80–90%) achieved by our method represents a significant improvement over these existing approaches.
Table 9: Comparison of feature lifting fidelity (Cosine Similarity). Red indicates the best performance (highest), Pink indicates the second best, and Green indicates the lowest performance. Our method consistently achieves the highest fidelity across all feature types.
Feature Method F. T R W. Mean
[…]
Figure 10: Feature Map Comparison: This figure compares our rendered features against the ground truth 2D features. The ground truth visualization is generated by applying PCA to the feature maps extracted directly from the input image by the foundation model. The rendered features represent the PCA of the features projected from our solved splats. These results demonstrate the SAM+OpenCLIP pipeline. Additionally, we compare our lifted result to current SOTA lifting method in Fig.13
Figure 11: Feature Map Comparison: This figure compares our rendered features against the ground truth 2D features. The ground truth visualization is generated by applying PCA to the feature maps extracted directly from the input image by the foundation model. The rendered features represent the PCA of the features projected from our solved splats. These results demonstrate the DINOv2 pipeline. Additionally, we compare our lifted result to current SOTA lifting method in Fig.14
Figure 12: Further Visualization results, we are using Teatime, ramen, figurines, waldo kitchen results from different scenes. More visualization results could be found in the supplementary materials
Figure 13: Qualitative Comparison of Feature Fidelity: We visualize the principal components (PCA) of the lifted feature fields rendered into 2D using the SAM+OpenCLIP backbone. We compare our proposed method (left columns) against another optimization-based Semantic Gaussian Guo et al. (2024) (middle column) and the heuristic Argmax Lifting Chacko et al. (2025) (right column). The input RGB and ground truth feature map are the same as the Fig.10. Our method recovers sharper object boundaries and more consistent internal feature representations (higher fidelity), avoiding the floating Gaussian artifacts and the noise inherent in heuristic aggregation.
Figure 14: Qualitative Comparison of Feature Fidelity: We visualize the principal components (PCA) of the lifted feature fields rendered into 2D using the DINOv2 backbone. We compare our proposed method (left columns) against another optimization-based Semantic Gaussian Guo et al. (2024) (middle column) and the heuristic Argmax Lifting Chacko et al. (2025) (right column). The input RGB and ground truth feature map are the same as the Fig.11. Our method recovers sharper object boundaries and more consistent internal feature representations (higher fidelity), avoiding the floating Gaussian artifacts and the noise inherent in heuristic aggregation.
For qualitative feature visualizations, please refer to Fig. 10, Fig. 11, and Fig. 12. Comparisons with existing feature-agnostic methods, such as Chacko et al. (2025); Guo et al. (2024), are provided in Fig. 13 and Fig. 14.
Kernel Agnosticism
For multiple kernel feature visualization, one can check Fig.8. There are plenty of the content in the appended video.
Appendix I Threshold